OpenStudy (ashleyisakitty):

What is the equation of the circle with center (3,5) that passes through the point (-4,10). Check my answer! :)

OpenStudy (ashleyisakitty):

i think its (x-3)^2+(y-5)^2=74

OpenStudy (agent47):

can you post the steps that got you there?

OpenStudy (ashleyisakitty):

no

OpenStudy (agent47):

looks good to me

OpenStudy (ashleyisakitty):

my two final choices are betwenn (x-(-4))^2+(y-10)^2=74 and my orinal choice. I am leaning with my first choice. Critic al urgent

OpenStudy (ashleyisakitty):

great

OpenStudy (agent47):

Don't guess though.. why does r^2 = 74?

OpenStudy (agent47):

darn it.. draw isn't working... hang on..

OpenStudy (ashleyisakitty):

It wasnt a guess

OpenStudy (agent47):

Sounded like an educated guess lol.. Anyway.. the equation of a circle is: (x-h)^2+(y-k)^2=r^2 where (h,k) is the center of the circle, and r is the radius.